MODS1 Commissioning Run 5: UTC 2011 February 23

Observer(s):J. Hill, A. Rakich (Tucson)
Telescope Operator: S. Allanson (LBTO)
Telescope Support: (LBTO)
Instrument Support: (LBTO)


Started active optics with MODS1 AGW up at midnight - TCS binocular testing with BP9 before that. Seeing 1-3" worst for first 2/3rds of night. Multiple failed acquisitions. Various bugs. THe field collect data not so good. The field check data better - Indicates large binodal astigmatism consistent with large M2 active optics offsets.

Binocular Testing of BP9

Monocular pointings to M5_0269

01:47 Adjust SX IE=-86 CA=+63 relative to the LUCI pointing model

01:58 Adjust DX IE=+17 CA=-4 relative to the IRTC pointing model

Binocular Offset Tests

Synchronized Preset Tests

02:14 Sync presets to BS9126 didn't get star on the left

02:xx Sync presets to M5_0269 ACQUIRE

Adjust SX IE=-92 CA=+62.5 relative to the LUCI pointing model

Adjust DX IE=+17 CA=-8 relative to the IRTC pointing model

02:xx Syn presets to BS9126

02:29 Async preset DX finds the star

02:41 Async ACTIVE SX

Binocular Guiding

02:50 Binocular guiding is working!
But active optics is broken on the right side. See IT 3193.

Somewhere here we switched to being LUCI on both sides.

PCS Locking Turned On

03:06 Sync presets BS9126 ACQUIRE
Star moved away, apparently because |Tip| changed from 30 to 20.

03:35 BOTH preset - gets star on SX, but not DX.
Adjust SX IE=-102 CA=+64 Tip=-36 Tilt=-13
Spiral DX, find star with horrible coma (needed -5000 Z8). Adjust DX IE=-38 CA=+28 wrong star?

Unlocked PCS

04:05 Restart IIF for patch, PCS locking off, GSR binning=2

04:xx Sync presets BS9143, but no stars

04:xx Sync presets ACQUIRE ACT0203
Adjust SX *IE=-115 CA=+60*
Adjust DX IE=-52 CA=+47

04:32 Sync presets ACTIVE BS9143 - GSCL crashed

04:50 stop/startAGW RFBG

Problem turned out to be that AIP_R GCS parameters for luci2 were old. See IT.

Unlocked we see both stars at SX Tip=-13.8 Tilt=-22.2

Locked PCS

JMH still has more notes to add here.

Mods active optics check out.

7:50 After a couple of failed presets, realized that no-one had updated GCS to version B5. Did that. Again failed acquisition a couple of times around 7:50 UT. Always failing at the third acquisition image. MODS is unusable while this problem persists.

7:55 Seeing 2-3 arc seconds.

7:55 had to restart idl due to failed acquisition. 1st off-axis measurements, coma small, z4, z5, z9 and z10 large. 2nd off-axis measurement; z4-z5, z10 largish 3rd off-axis measurement, did it find the right guide star?, this one looks faint. Low aberration numbers but suspect the star was too faint. Lots of focus when we went back on-axis.
8:12 UT, FOR some reason, only did three cycles on axis collimating had a WFE of 500 nm then went off axis to measure?????? 4th off-axis measurement, lowish aberration, Z5 and Z10 stand out a bit. 2.5" seeing.
8:17 Again, three collimations and it took of to measure even though it was definitely not converged and there was no prompt window to keep collimating. This is going to complicate data reduction significantly. 6th off axis guide star, coma looks great, z5, z6 z9 and z10 don't. Could just be the conditions.
8:30 had another fail acquiring AGW image causing IDL to crash (happend a lot last week), started again with START=7. 8:37 failed preset, resending with START = 8.

8:47 Failed acquisition step 3 again. John changing Sextractor flag (sextractorFlag) back to 16 (we tried 20 and got this new type of failure).

8:49 resending preset START=7. Acquisition failed again. New type of failure continues.

8:50 try different PA - maybe this PA is unfavourable for stray light... START=10.

9:15 Note the collimation prior to presetting to GS 14. collimated 3 times. Did not converge to 300 nm. No prompts. Went off to measure anyway.

9:31 very last on-axis collimation preset failed, resent preset but there is an extra delay here.

Field Check on BS9169

9:35 now doing field_check on BS9169, starting a little low (but better than where we just were) but this is the only target away from wind and moon.

9:40. Numerous Information file not found errors.

9:44 new acquisition error.

9:45. This info file error is very bothersome and is happening all the time now. In this instance it had the WFSC file, reported a WFE then said it couldn't find the info file

9:46 John resetting his flag (affecting acquisition) to 14.

9:51 a couple of failed presets when we try to go to the first star on the list with the following error message:

Invalid guide probe Y position '137.616', must be -204..2 

I'm guessing that this error is in Doug's specified PA. Trying field_check /GCS, START=2.

9:54 New error:
===> Error in irc_presetTelescope
        err = Resmsg[0]  = PresetTelescope result status: Error
        err = Resmsg[1]  = WARNING: guide cam exposure time is less than 2000 ms which might cause guiding on atmospheric dispersions
        err = Resmsg[2]  = adjusted exposure time to 100 ms
        err = Resmsg[3]  = error: probe offset movement drove star off the CCD.

John changing his parameter back to 16 and trying again.

10:00 Every time we try to go to an off-axis star with field check we get this (when all of the other errors stay away):
 => Collimated in 2 iterations
      # Target star: 0   BS9169
        17:13:44.9407 +54:33:17.960 RADEC 2000.0 J2000 0.0 0.0 11.64 U 0.78 B_V 0.3
      # Guide star: 23
        17:13:14.8931 +54:31:06.461 RADEC 2000.0 J2000 0.0 0.0 12.45 U 0.88 B_V 0.5
      # Y-Axis Position Angle:   63.35
        request = irc PresetTelescope -116.65 POSITION ACTIVE left; echo $?
===> Error in irc_presetTelescope
        err = Resmsg[0]  = PresetTelescope result status: Error
        err = Resmsg[1]  = AGw command status:  Invalid guide probe Y position '173.866', must be -204..2
        err = Resmsg[2]  = error: setProbePosition: couldn't move probe to position (SFP) (x,y,z): 0.74304, 173.967, 0 mm.
        err = Resmsg[3]  = guidestar list successfully changed but probe couldn't be moved to position.
% Program caused arithmetic error: Floating illegal operand

10:05. John may have found the problem, an occurance of "MODS1" in the field_check code that should have been "MODS". This was making the PAs wrong by 180.

10:20 Seems I sent "field collect" instead of "field_check" after John's change, resending field_check on "BS9169_short".

10:25 first acquisition failed, sent again, worked.

First field check off-axis, PA25 largish focus and Z5.
10:31 failed preset on the first back-to-on-axis-to-collimate step. Flag on identified star was 24 in this case hence the failure. Trying again... PA25 largish focus z5 (again). going back on axis
again, three iterations and we're off even though the last WFE was 420 nm.
10:47. Failed preset, this time illegal commanded probe position, MODS probably can't accomodate +/- 60 degrees PA, hardwiring for 40 degrees and trying again.
10:50. Now have "" to try. Restarting idl and trying again...

10:51 large active optics offset from current model (including current globals) of -1.781 mm Y 186" Rx).

First field check off-axis, PA25 largish focus and Z5. This is a consistent pattern at least...
back on axis to collimate: PA65 smaller focus larger z5, z8 and z9 are both a litte larger too... Still, FAR LESS variation than last time we did this, and in significantly worse seeing. PA-15 largish z4, z5, everything else small. Great. So the Z4 and Z5 seem to be consistent features at different PAs (unlike before). on axis to collimate in one iteration, now off to next star.

Second field check off-axis:

PA64 Z4 dominates. Large positive value. (700nm) back on-axis AZcam / GCS error.

11:19 Try field_collect_40 again with START=2.

Same error. Not getting WFSC images. Stopped GCS.

11:24 cycled power on WFS camera (we think) and try again with START=2...

11:25. Problem persists.

11:40 Hand the telescope to the power guys 20 minutes early.

11:50 AzCam server says "readout counter stuck". JMH cured the problem by rebooting the AzCam PC over VNC.

12:00 MODS stuff powered off in preparation for the outage.

-- AndrewRakich - 23 Feb 2011
Topic revision: r5 - 25 Feb 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback